# Tidally Delayed Spin-Down of Very Low Mass Stars

### v2

@@ 1 Introduction
The observed period distributions of main sequence stars as a function of age constrain models of stellar spin evolution. In turn, understanding the spin evolution is important for gyrochronological aging (Barnes, 2003; Reiners and Mohanty, 2012), which depends on a one-to-one mapping between rotation period and stellar age bolstered by a predictive theory. However, observations reveal that the rotation and activity of significant populations of main sequence stars evolve anomalously with age (Meibom et al., 2009, 2011; Agüeros et al., 2018), challenging standard theoretical models restricted to the magnetically mediated spin down of isolated stars (Matt et al., 2012; Reiners and Mohanty, 2012; van Saders and Pinsonneault, 2013; Gallet and Bouvier, 2013; Matt et al., 2015; van Saders et al., 2016). (Newton et al., 2016). Observations from multiple surveys (Irwin et al., 2011; Newton et al., 2016; Kado-Fong et al., 2016; Newton et al., 2018; Howard et al., 2020; Popinchalk et al., 2021; Shan et al., 2024) consistently reveal a bimodal distribution in rotation period for M dwarfs with a notable dearth of late-type M dwarfs with intermediate rotation periods. This bimodality is characterized by populations of rapid and slow rotators ($>100$ days) with a significant lack of intermediate rotators between roughly 10-40 days (Kado-Fong et al., 2016) and in some cases up to 50 (Shan et al., 2024) or 70 days (Newton et al., 2016). This gap can be interpreted as evidence of a rapid spin-down of M dwarfs.
@@
While a change in magnetic topology might help to explain some anomalies (Metcalfe et al., 2016; Garraffo et al., 2018), many stars have stellar, substellar, or planetary companions whose influence on spin angular momentum evolution must also be understood. M dwarfs with short rotation periods may either be very young, or involve short-period stellar binaries (McQuillan et al., 2013). Using galactic kinematics to estimate M dwarf ages, there is in fact a sub-population of fast-rotating old stars ($7-13$ Gyr) (Irwin et al., 2011). The presence of a sufficiently massive companion can modify the spin evolution of the host stars via equilibrium and dynamical tides (Zahn, 1977; Ogilvie and Lin, 2007; Ogilvie, 2013; Mathis, 2015) and we shall see that this can generate the gap in the rotation period distribution. Even sub-stellar brown dwarf (BD) companions may be sufficient to explain this trend for the sub-population of old, late-type M stars.
@@
In contrast to semi-empirical (Matt et al., 2015) or semi-analytical and numerical magnetohydrodynamic models (Gallet and Bouvier, 2013, 2015; Amard et al., 2016; Garraffo et al., 2018), our model herein is derived from first principles, starting with a dynamical MHD approach (Blackman and Owen, 2016) that is augmented to include tides. While existing models typically use parametrized torque laws based on MHD simulations or empirically tuned to match observations, they often treat stellar properties as independent or use external prescriptions and empirical scaling laws to couple them. Our model dynamically and self-consistently couples the evolution of X-ray luminosity, rotation, mass loss, magnetic field strength to tidal effects of the companion and its orbital evolution. Previous work has also focused on solar-type stars and the effects of core-envelope decoupling in studying the influence of a massive planet on the stellar surface rotation (Gallet and Bouvier, 2013, 2015). Here we study the influence of tides on the stellar spin evolution of fully convective M dwarf-BD systems for a range of companion masses and orbital separations. We focus on explaining the dearth of intermediate rotators for late-type M dwarfs (Newton et al., 2016) by accounting for equilibrium tides. For comparison, we also investigate the spin evolution implications of sub-BD mass companions for partially convective K dwarfs and early M dwarfs.
@@ 2.1 Magnetic spin-down model for an isolated magnetized star
In our theoretical spin evolution model that dynamically couples stellar spin, magnetic field strength, X-ray luminosity and mass loss, dynamo-generated magnetic fields are the main coronal energy source, supplying both open and closed coronal fields. Hot gas propagates along open field lines and removes angular momentum, while magnetic dissipation in closed field lines sources X-ray emission. More detailed derivations are in (Blackman and Owen, 2016) where the original model was applied to the Sun, and it was further developed in (Kotorashvili et al., 2023) and applied to older-than-solar low-mass MS stars. The main ingredients are an isothermal Parker wind (Parker, 1958); angular momentum loss from the equatorial plane (Weber and Davis, 1967); a dynamo saturation prescription based on a combination of magnetic helicity evolution and buoyant loss of magnetic fields (Blackman and Field, 2002; Blackman and Brandenburg, 2003), used to estimate the magnetic field strength; replacement of the convection time in the dynamo coefficients by shear or rotation time for fast rotators (Blackman and Thomas, 2015); and a coronal equilibrium condition (Hearn, 1975) that determines how the magnetic energy supply is distribiuted into radiation, wind, and conductive losses. Together, these components lead to a system of coupled equations for the X-ray luminosity, magnetic field, spin evolution and mass loss below.
The first necessary equation is the relation between X-ray luminosity $l_{\rm x{\star}}(t)$ and radial magnetic field $b_{r{\star}}(t)$:
@@
Compared to the original model of (Blackman and Owen, 2016), here we have generalized the prescription for the correlation timescale. Instead of using a shear parameter111 $|\Omega_{0}-\Omega(r_{c},\theta_{s})|=\Omega_{0}/s$, where $\Omega$ is surface rotational speed; $\theta_{s}$ is a fiducial polar angle; $r_{c}$ is a fiducial radius in the convective zone. $s$ we use the error function $\rm erf(\epsilon/\tilde{R}o_{\star})$ which depends on the parameter $\epsilon$ and Rossby number $\tilde{R}o_{\star}$. Here $\lambda$ is the exponent of the power law dependence of magnetic starspot area covering fraction on X-ray luminosity (Nichols-Fleming and Blackman, 2020). Here $g_{L}(t)=({1.4-0.4t})^{-\frac{\lambda-1}{4}}$ measures the bolometric luminosity evolution on the main sequence. For simplicity, we take $g_{L}(t)=1$ for both partially and fully convective stars. More details on the numerical values of parameters used are in section 2.4.
Next, we have the equation for the toroidal magnetic field:
@@
where $b_{\phi{\star}}(t)$ is the dimensionless toroidal magnetic field, normalized to the initial toroidal magnetic field $B_{\phi,*i}$  [^2] We substitute subscript $\rm i$ for $\rm n$ used in (Blackman and Owen, 2016), as here we use initial stellar values for rather than present-day values.. Quantities $\dot{M}_{*i}$, $B_{\rm r,\star i}$, $\Omega_{\star i}$ are initial mass loss, radial magnetic field, and angular velocity, respectively; $\dot{m}_{\star}(t)$ is a mass loss rate (see equations 4 and 5 below); $M_{\star}$ and $R_{\star}$ are the stellar mass and radius, assumed to be constant; the Alfvén radius $r_{A}$ is given below in equation (6); $\omega_{\star}(t)=\Omega_{\star}(t)/\Omega_{\star i}$ is the dimensionless angular velocity described by
@@
where $t_{\star}$ is the normalization parameter taken as $1$ Gyr. Here $\beta_{\star}$ is the radius of gyration; $q$ is a dimensionless moment of inertia parameter that depends on the assumed constant fraction of the star in which the field is anchored; $\tilde{u}_{A}(t)$ and $u_{A,{\star}i}$ are the dimensionless Alfvén speed and initial value for Alfvén speed; and $m_{\star}=1$ is the dimensionless stellar mass in units of initial stellar mass, which does not evolve.
The fourth equation needed for the model is the mass loss rate, for which we have two regimes. For the thermal conduction loss dominated regime - Regime I:
@@
where $\tilde{T}_{0\star}(t)$ is a dimensionless coronal temperature at equilibrium, normalized such that transition between regime I and regime II occurs at $\tilde{T}_{0\star}=0.5$. The $\tilde{T}_{0,{\star}i}$ represents the initial coronal temperature.  [^3] Here we have included the correction (Kotorashvili et al., 2024) to (Kotorashvili et al., 2023). Here $m_{\odot\star}=\frac{M_{\star}}{M_{\odot}}$ and $r_{\odot\star}=\frac{R_{0}}{R_{\odot}}\sim\frac{R_{\star}}{R_{\odot}}$, where $R_{0}$ represents radius at the coronal base; $M_{\odot}$ and $R_{\odot}$ are solar mass and radius respectively.
The Alfvén radius needed for equation (2) and derived following (Weber and Davis, 1967; Blackman and Owen, 2016; Lamers and Cassinelli, 1999) is
where the Alfvén speed is given by
@@
where $W_{h}[-D(r_{A})]$ is the Lambert W function for the Parker wind with $h=0$ for $r\leq r_{s}$ and $h=-1$ for $r\geq r_{s}$ (Cranmer, 2004), and
@@ 2.2 Influence of tides on the evolution of stellar spin
The spin evolution model summarized in the previous section must be generalized to account for tidal stresses, which can exchange angular momentum between the orbit and the spins of the primary and secondary when the primary star has a close stellar or sub-stellar companion. We assume that tidal stresses are redistributed much faster than the secular spin evolution that we will solve for (Goupil and Zahn, 2008; Rozelot and Neiner, 2013). We focus only on equilibrium tides, for which the star evolves in hydrostatic equilibrium during its response to tidal forcing (e.g, (Zahn, 1977; Tassoul, 2000)).
We consider a sub-stellar companion in a circular orbit, and a uniform rotation of the primary star at angular velocity $\Omega_{\star}$. The gravitational field of the secondary deforms the primary, and vice-versa, but we focus on the former. Since the star is not a perfect fluid, it will incur a viscously mediated time lag $\Delta t$ to the deformation that offsets the direction of tidal bulges from the instantaneous direction of the external force by a small angle $\delta$, producing a torque. In the weak-friction approximation $\delta\ll 1$, the tidal torque $\Gamma_{\rm T}$ on the primary can be written (e.g (Tassoul, 2000)):
@@
where $m$ is the secondary (companion) mass and $a$ is the orbital separation. Here $k$ is the apsidal motion constant determined by the structure of the star.
@@
The time $t_{\rm conv}$ is a characteristic timescale for kinetic energy dissipation in the stellar convection zone governed by turbulent viscosity and can be written as $t_{\rm conv}=R_{\star}^{2}/\nu_{t}\simeq\left(M_{\star}R_{\star}^{2}/L_{\star}\right)^{1/3}$ , where $\nu_{t}$ is eddy viscosity and $L_{\star}$ is the stellar bolometric luminosity (Zahn, 1989). This is a crude but useful order-of-magnitude scaling for studying tidal dissipation in convective envelope.

### v3

@@ 1 Introduction
The observed period distributions of main sequence stars as a function of age constrain models of stellar spin evolution. In turn, understanding the spin evolution is important for gyrochronological aging (barnes2003rotational; Reiners2012), which depends on a one-to-one mapping between rotation period and stellar age bolstered by a predictive theory. However, observations reveal that the rotation and activity of significant populations of main sequence stars evolve anomalously with age (meibom2009stellar; meibom2011kepler; agueros2018new), challenging standard theoretical models restricted to the magnetically mediated spin down of isolated stars (Matt2012; Reiners2012; vanSaders2013; Gallet2013improved; Matt2015; vanSaders2016). (Newton2016). Observations from multiple surveys (irwin2011Angular; Newton2016; Kado-Fong2016; Newton2018; Howard2020; popinchalk2021evaluating; Shan2024carmenes) consistently reveal a bimodal distribution in rotation period for M dwarfs with a notable dearth of late-type M dwarfs with intermediate rotation periods. This bimodality is characterized by populations of rapid and slow rotators (>100>100 days) with a significant lack of intermediate rotators between roughly 10-40 days (Kado-Fong2016) and in some cases up to 50 (Shan2024carmenes) or 70 days (Newton2016). This gap can be interpreted as evidence of a rapid spin-down of M dwarfs.
@@
While a change in magnetic topology might help to explain some anomalies (Metcalfe2016; Garraffo2018), many stars have stellar, substellar, or planetary companions whose influence on spin angular momentum evolution must also be understood. M dwarfs with short rotation periods may either be very young, or involve short-period stellar binaries (McQuillan2013). Using galactic kinematics to estimate M dwarf ages, there is in fact a sub-population of fast-rotating old stars (7137-13 Gyr) (irwin2011Angular). The presence of a sufficiently massive companion can modify the spin evolution of the host stars via equilibrium and dynamical tides (Zahn1977; Ogilvie+lin2007; Ogilvie2013; Mathis2015b; benbakoura2019evolution; ahuir2021dynamical; Pezzotti2025) and we shall see that this can generate the gap in the rotation period distribution. Even sub-stellar brown dwarf (BD) companions may be sufficient to explain this trend for the sub-population of old, late-type M stars.
@@
In contrast to semi-empirical (Matt2015) or semi-analytical and numerical magnetohydrodynamic models (Gallet2013; Gallet+Bouvier2015; Amard2016; Garraffo2018; benbakoura2019evolution; ahuir2021dynamical; Pezzotti2025), our model herein is derived from first principles, starting with a dynamical MHD approach (Blackman+owen2016) that is augmented to include tides. While existing models typically use parametrized torque laws based on MHD simulations or empirically tuned to match observations, they often treat stellar properties as independent or use external prescriptions and empirical scaling laws to couple them. Our model dynamically and self-consistently couples the evolution of X-ray luminosity, rotation, mass loss, magnetic field strength to tidal effects of the companion and its orbital evolution. Previous work has also focused on solar-type stars, the effects of core-envelope decoupling and dynamical tides in addition to equilibrium tides in studying the influence of a companion on the stellar surface rotation and X-ray activity (Gallet2013; Gallet+Bouvier2015; Amard2016; Rao2021; benbakoura2019evolution; ahuir2021dynamical; Pezzotti2025). Here we study the influence of tides on the stellar spin evolution of fully convective M dwarf-BD systems for a range of companion masses and orbital separations using a simple prescription for tidal torque. We focus on explaining the dearth of intermediate rotators for late-type M dwarfs (Newton2016) by accounting for equilibrium tides. For comparison, we also investigate the spin evolution implications of sub-BD mass companions for partially convective K dwarfs and early M dwarfs.
@@ 2.1 Magnetic spin-down model for an isolated magnetized star
Our spin evolution model that dynamically couples stellar spin, magnetic field strength, X-ray luminosity and mass loss is based on the analytic framework of (Blackman+owen2016), where the original model was derived and applied to the Sun, and was further developed in (kbo2023) for application to older-than-solar low-mass MS stars. In this conceptual framework, dynamo-generated magnetic fields are the main coronal energy source, supplying both open and closed coronal fields of strength proportional to each other. Hot gas propagates along open field lines and removes angular momentum, while X-ray emission from magnetic dissipation results from reconnecting closed field lines that can subsequently open up. The main dynamically coupled ingredients are: an isothermal Parker wind (Parker1958); angular momentum loss from the equatorial plane (Weber1967); a dynamo saturation prescription based on a combination of magnetic helicity evolution and buoyant loss of magnetic fields (Blackman2002; Blackman2003), used to estimate the magnetic field strength; prescription for the turbulent correlation time in the dynamo coefficients that asymptotes to the convection time for slow rotators and shear or rotation time for very fast rotators (Blackman2015); and a coronal equilibrium condition (Hearn1975) that determines how the magnetic energy supply is divided into radiation, wind, and conductive losses. Together, these components lead to a system of coupled equations for the X-ray luminosity, magnetic field, spin evolution and mass loss below. More detailed derivations are in (Blackman+owen2016) and (kbo2023).
The first key equation that emerges from the model is the relation between dimensionless X-ray luminosity $l_{\rm x{\star}}(t)$ and radial magnetic field $b_{r{\star}}(t)$:
@@
This surface radial magnetic field originates from dynamo-produced toroidal interior field that emerges to the surface. The $b_{r}$ is interpreted as the azimuthally and dynamo cycle-averaged poloidal field which we scale as local radial field.The associated X-ray luminosity is a product of the magnetic energy flux averaged over a dynamo cycle times the solid angle fraction $\Theta/4\pi$ through which magnetic field rises, which approximately equals the areal fraction of sunspots during the cycle. Here $\lambda$ is the exponent of the power law dependence of the magnetic starspot area covering fraction $\Theta$ on X-ray luminosity $\mathcal{L}_{X}$, namely $\Theta\propto\mathcal{L}_{X}^{\lambda}$. We take $\lambda=1/3$, which is consistent with the range inferred from observations of star spot covering fractions for FGKM stars (NicholsFleming2020). The field that emerges over a surface area consistent with the sunspot covering fraction $\Theta$ is $|B_{r\star}|=\Theta b_{r\star}\simeq\mathcal{L}_{X}^{\lambda}b_{r\star}\propto l_{\rm x{\star}}^{\lambda}b_{r\star}\propto b_{r{\star}}^{(\frac{4}{1-\lambda})\lambda}b_{r\star}$. For $\lambda=1/3$ we have $|B_{r\star}|\propto b_{r{\star}}^{3}$ and $l_{\rm x{\star}}\propto|B_{r\star}|^{2}.$, which is in a good agreement with observed relationship between magnetic flux and X-ray luminosity (Pevtsov2003; Vidotto2014; See2015; Zhuleku2020).
[…]
Next, we consider angular momentum loss from equatorial plane following (Weber1967; Lamers1999) to find the azimuthal component of magnetic field at the base of the corona which we refer to as the ”coronal toroidal field” and the equation for angular velocity. We assume a Parker spherical wind (Parker1958) which propagates along radial large-scale fields through the Alfvén radius $r_{A}(t)$, which following (Weber1967; Blackman+owen2016; Lamers1999) is given by,
$\frac{r_{A}(t)}{R_{\star}}=\left(1-\frac{R_{\star}B_{r,\star i}B_{\phi,\star i}}{\dot{M_{\star i}}\Omega_{\star i}}\right)^{1/2}=\left(1+\frac{R_{\star}|B_{r,\star i}||B_{\phi,\star i}|}{\dot{M_{\star i}}\Omega_{\star i}}\right)^{1/2},$ (2)
where quantities $\dot{M}_{\star i}$, $B_{\rm r,\star i}$, $B_{\rm\phi,\star i}$, $\Omega_{\star i}$ are the initial mass loss, radial magnetic field at the coronal base, coronal toroidal magnetic field, and angular velocity, respectively. Here $R_{\star}$ is the stellar radius and assumed to be constant since we are focusing on stars already on the main sequence.
Combining equation (2) with a separate expression for $r_{A}(t)$ that is independent of coronal toroidal magnetic field, we obtain the coronal toroidal magnetic field equation. As derived from the mass loss rate to outflow speed relation, the definition of the Alfvén radius, and the radial field fall off of $1/r^{2}$, $r_{A}(t)$ is given by
where $\dot{m}_{\star}(t)$ is the mass loss rate (see equations 7 and 8 below); $\tilde{u}_{A}(t)$ and $u_{A,{\star}i}$ are the dimensionless Alfvén speed and initial value for Alfvén speed given in equation (9) below. Combining equations (2) and (3) gives
@@
where $b_{\phi{\star}}(t)$ is the dimensionless coronal toroidal magnetic field, normalized to the initial coronal toroidal magnetic field $B_{\phi,\star i}$  [^2] We substitute subscript $\rm i$ for $\rm n$ used in (Blackman+owen2016), as here we use initial stellar values for rather than present-day values.. $M_{\star}$ is the stellar mass and assumed to be constant; $\omega_{\star}(t)=\Omega_{\star}(t)/\Omega_{\star i}$ is the dimensionless angular velocity. From conservation of angular momentum in the Weber1967 framework, for which the magnetic field is nearly rigid out to the Alfvén radius, the stellar angular velocity evolves as
$\frac{d\Omega_{\star}(t)}{dt}=-\frac{q\dot{M}_{\star}(t)\Omega_{\star}(t)}{\beta^{2}M_{\star}}\frac{r_{A}(t)}{r_{\star}}^{2}$ (5)
and dimensionless angular velocity is described by
@@
where $t_{\star}$ is the normalization parameter taken as $1$ Gyr. $m_{\star}=1$ is the dimensionless stellar mass in units of initial stellar mass, which does not evolve. Here $q$ is a dimensionless moment of inertia parameter that depends on the assumed constant fraction of the star in which the field is anchored. We use $q=1$ for all stars, which is equivalent to assuming that the full stellar mass is coupled to the anchoring magnetic field that contributes to spin-down; $\beta_{\star}$ is the radius of gyration taken from stellar evolutionary models computed using the Modules for Experiments in Stellar Astrophysics package (MESA) (Claret2023). For a stellar mass $M_{\star}=0.2M_{\odot}$ we use $\beta_{\star}=0.458$, while for $M_{\star}=0.6M_{\odot}$ we take $\beta_{\star}=0.359$.
Time dependent data are needed to accurately assess the consistency of equation (4) but we can get an approximate time averaged prediction by using our numerical solutions to find the scaling between toroidal and observed radial magnetic fields $b_{\phi}(t)=Cb_{r}(t)^{\zeta(t)}\propto C|B_{r}(t)|^{\zeta(t)/3}$ where the proportionality follows from the discussion below equation (1), and we infer $\zeta(t)$ by solving for $b_{\phi}(t)$ as a function of $b_{r}(t)$. If we approximate $\zeta(t)$ from $t\approx 1.5-10$ Gyr by a constant for the systems considered in our paper, we find that $C$ - $10^{-1},10^{-3}]$ and $\eta=\zeta/3$ falls into the range between $[0.7,1]$, roughly comparable to the observed scalings between toroidal field energy and poloidal field energy. (See2015). More detail on our estimate for $\eta$ is given in the supplement Fig. 6.
In this model, dynamo generated magnetic fields source the energy for the wind and X-rays or are dissipated through thermal conduction. To determine the dominant flow path of magnetic energy we use the coronal flux equilibrium condition (Hearn1975) between wind flux, conductive loss and the radiative (X-ray) loss, assuming that the equilibrium is established over shorter time-scales compared to the secular gigayear evolution time. For detailed derivations refer to Blackman+owen2016 and kbo2023 where we find the relation between mass loss and X-ray for two regimes: regime I as the lifetime phase of a star for which thermal conduction flux dominates the outflow flux and regime II when the reverse is true. For the thermal conduction loss dominated regime - Regime I:
@@
where $\tilde{T}_{0\star}(t)$ is a dimensionless coronal temperature at equilibrium, normalized such that transition between regime I and regime II occurs at $\tilde{T}_{0\star}=0.5$. The $\tilde{T}_{0,{\star}i}$ represents the initial coronal temperature.  [^3] Here we have included the correction (kbo2023errata) to (kbo2023). Here $m_{\odot\star}=\frac{M_{\star}}{M_{\odot}}$ and $r_{\odot\star}=\frac{R_{0}}{R_{\odot}}\sim\frac{R_{\star}}{R_{\odot}}$, where $R_{0}$ represents radius at the coronal base; $M_{\odot}$ and $R_{\odot}$ are solar mass and radius respectively. Blackman+owen2016 studied the wind dominated regime for the Sun, and found that equation (8) is consistent with the data.
Alfvén speed is given by
@@
where $W_{h}[-D(r_{A})]$ is the Lambert W function for the Parker wind (Parker1958) with $h=0$ for $r\leq r_{s}$ and h=1h=-1 for $r\geq r_{s}$ (Cranmer2004), and
@@ 2.2 Influence of tides on the evolution of stellar spin
The spin evolution model summarized in the previous section must be generalized to account for tidal stresses, which can exchange angular momentum between the orbit and the spins of the primary and secondary when the primary star has a close stellar or sub-stellar companion. We assume that tidal stresses are redistributed much faster than the secular spin evolution that we will solve for (Zahn2008; bookrozelot2012environments). We focus only on equilibrium tides, for which the star evolves in hydrostatic equilibrium during its response to tidal forcing. Dynamical tides can enhance the tidal influence but compared to equilibrium tides they become relevant only for a short time fraction of the overall evolution. Including these effects can be a subject for future work. (e.g., Zahn1977; tassoul_2000).
We consider a sub-stellar companion in a circular orbit, and a uniform rotation of the primary star at angular velocity $\Omega_{\star}$. The gravitational field of the secondary deforms the primary, and vice-versa, but we focus on the former. Since the star is not a perfect fluid, it will incur a viscously mediated time lag $\Delta t$ to the deformation that offsets the direction of tidal bulges from the instantaneous direction of the external force by a small angle $\delta$, producing a torque. We adopt the weak friction approximation. In the weak-friction approximation $\delta\ll 1$, assuming that the tidal force acts as a perturbation on the primary. Then the tidal torque $\Gamma_{\rm T}$ on the primary can be written (e.g., tassoul_2000):
@@
where $m$ is the secondary (companion) mass and $a$ is the orbital separation. Here $k$ is the apsidal motion constant determined by the structure of the star, where we use values $\rm k=0.155$ and $\rm k=0.054$ from MESA (Claret2023) corresponding to the stellar masses $M_{\star}=0.2M_{\odot}$ and $M_{\star}=0.6M_{\odot}$.
@@
The time $t_{\rm conv}$ is a characteristic timescale for kinetic energy dissipation in the stellar convection zone governed by turbulent viscosity and can be written as $t_{\rm conv}=R_{\star}^{2}/\nu_{t}\simeq\left(M_{\star}R_{\star}^{2}/L_{\star}\right)^{1/3}$ , where $\nu_{t}$ is eddy viscosity and $L_{\star}$ is the stellar bolometric luminosity (Zahn1989). This is a crude but useful order-of-magnitude scaling for studying tidal dissipation in convective envelope.